OpenStudy (anonymous):

( am i right? I was thinking it was b.) What is the volume of a ball if the diameter if the ball is 6 inches? Leave your answer in terms of π. A) V = 12 π cubic inches B) V = 36 π cubic inches C) V = 288 π cubic inches D) V = 27.75 π cubic inches

OpenStudy (paki):

hmm.. simple... just use the formula... and solve... volume of a ball= V = (4/3) * pi * radius^3... put the values and get your answer.... @Heypeoplz

OpenStudy (anonymous):

ok so its A. 12 pie

OpenStudy (paki):

no... :-)

OpenStudy (paki):

what value you used for radius...?

OpenStudy (anonymous):

isnt it 3

OpenStudy (anonymous):

Is this the equation i'm suppose to do (4/3)pi(3)^3

OpenStudy (paki):

yup 3 is rite... then you have some mistake in calculation...

OpenStudy (anonymous):

this is what i am suppose to solve right? (4/3)pi(3)^3

OpenStudy (paki):

V = (4/3) * pi * radius^3 V= (4/3) * pi * (3)^3 V= ...?

OpenStudy (anonymous):

I got 36 pi

OpenStudy (paki):

yup solve now, then tell me... take care of the power of "3"

OpenStudy (paki):

ahhhhh... yes, now you are rite :-)

OpenStudy (anonymous):

So i was right, b is the answer.

OpenStudy (paki):

yah

OpenStudy (anonymous):

thanks :)

OpenStudy (paki):

pleasure :-)